# Code Graph Model (CGM): A Graph-Integrated Large Language Model for Repository-Level Software Engineering Tasks

### v2

@@ C.1.1 Training Data
As mentioned in section 4, we construct training data for different training phase of CGM respectively. Meanwhile, to enhance the model’s ability in code completion, we also construct code completion samples for fine-tuning of code completion task. To explore the performance of CGM across different programming languages, our data includes both Python and Java. When constructing the training data, we filter out the repositories involved in the test sets for testing to avoid data leakage. Data for Subgraph Reconstruction Pre-training: We obtain 500k Python and 360k Java subgraphs (with the maximum length of 8k tokens) from a total of 20k high-star Github repositories. Data for Issue Fixing: We collect 200k issue-patch pairs (100k per language), from GitHub pull-requests. Among the 100k Python pairs, 14k are sourced from the SWEBench training set [8]. Data for Code Completion: The code completion samples are self-constructed from the above repositories, 250k per language.
@@ C.1.2 CGM-Multi
We initialize CGM-Multi with Qwen2.5-72B-Instruct [20] as the base LLM. Then pre-train it using subgraph reconstruction data and fine-tuning data (issue-fixing and code completion) in both two languages (Python and Java). To ensure balance between different languages, we use 360k subgraph reconstruction data for both languages. Training uses 4-bit QLoRA on 64 A100s (batch=32, lr=1e-4, and epoch=2). The encoder combines CodeT5+ [51] with LoRA (rank=64, and alpha=32), and the adapter uses a two-layer MLP with GELU activation. The first layer of the adapter maps the CodeT5+ output dimension of 256 to 8192, and the second layer maintains the dimension of 8192, which aligns with the LLM’s hidden dimension. We adopt Xformers [68] for efficient attention computation.
@@ C.5.2 Datasets
The following three benchmarks, which focus on repository-level issue fixing, are all evaluated using the Docker executable environment. SWE-bench Lite [5]: SWE-bench Lite contains 300 self-contained repository-level issues from 11 repositories, designed to test the model’s understanding of repository-level code changes and its ability to generate correct patches primarily focused on Python. It provides a realistic software engineering environment, including execution contexts, to evaluate the model’s ability to resolve real-world issues. SWE-bench Verified [69]: SWE-bench Verified contains 500 self-contained repository-level issues from 12 repositories. This dataset contains samples that have been verified to be non-problematic by human annotators. SWE-bench-java Verified [6]: This dataset include 91 Java issues from 6 repositories, enabling cross-language evaluation. Like SWE-bench Lite, it provides execution environments to validate the correctness of generated patches.
@@ C.6.2 Datasets
CrossCodeEval [62] is an emerging benchmark for cross-file code completion, which is constructed from a wide range of real-world repositories from GitHub in four popular programming languages: Python, Java, TypeScript, and C#. In our experiments, we evaluate the model’s code completion ability on only two languages, Java and Python. As shown in Table 9, we provide the dataset statistics of the CrossCodeEval benchmark for Java and Python. ComplexCodeEval [70] is a new benchmark for evaluating the performance of large code models in complex development scenarios. It includes 3,897 Java samples from 1,055 Java code repositories and 7,184 Python samples from 2,107 Python code repositories. Following the original setup of this benchmark, we randomly selected 100 samples each in Python and Java for the evaluation. Table LABEL:tab:_complex_python and Table LABEL:tab:_complex_java show the information of the selected samples. When evaluating the code completion capability, unlike the original setup which requires the model to complete the second half of the function, we ask the model to complete the middle line of the function given the contextual information. Table 9: The statistics of the CrossCodeEval for Java and Python. Feature Java Python # Repositories 239 471 # Files 745 1368 # Examples 2139 2665
@@ C.6.4 Baselines: Base Model
Given the subgraph same to CGM (see Appendix C.6.1), we textualize the graph into text sequnce and input into the following baselines based on the prompt templates provided by each benchmark [62, 70]. Since these base models have limitations on context length, we perform truncation on text inputs larger than 8k. Mistral-Large-2 [71] is a model developed by Mistral AI with 123 billion parameters. It stands out with a remarkable 128k tokens context length, proficiently handling dozens of languages and over 80 programming languages, excelling in code generation, mathematics, and reasoning. To be specific, we chose Mistral-Large-Instruct-2411 as the latest version of Mistral-Large to compare with CGM. DeepSeek-V2.5 [14] is a strong Mixture-of-Experts (MoE) language model characterized by economical training and efficient inference. It comprises 236B total parameters, of which 21B are activated for each token. Qwen2.5 [20] is a decoder-only LM series whose size varies from 0.5B to 72B trained on 18 trillion tokens. Its context length support up to 128K tokens and can generate up to 8K tokens. Qwen2.5-Coder [25] is a series of code-specific language models developed by Alibaba. Derived from Qwen2.5, it comes in six sizes, is trained on a vast 5.5-trillion-token corpus, and excels in various code-related tasks, outperforming many models of the same or larger size. Since it uses a specific format and tokens for training on the fill-in-middle code completion task, we followed this prompt setting during the evaluation to obtain its true performance. For the inference of the baseline model, we all deployed the above models using the VLLM framework with the model‘s default settings. All models inference on 4 A100s with 80G VRAM, except for DeepSeek-V2.5, which requires 8 A100s.
@@ C.6.5 Baselines: RAG Method
BM25 [72] is a classic information-retrieval algorithm based on the probabilistic model. Its core idea is to rank documents based on the relevance between query and documents. It serves as a traditional retrieval method that does not regard the structural information naturally existing in the coding task, and only performs similarity matching based on word frequency and text length. It was used in the original CrossCodeEval dataset to search for cross-file information based on the code snippets. In our experiments, we directly use the BM25 results provided by CrossCodeEval. R2C2-Coder [65] is a method that aims to enhance and benchmark the real-world repository-level code completion abilities of code Large Language Models. In particular, $R^{2}C^{2}$-Enhance reduces cross-file information to skeleton [^3] The file skeleton is a hierarchical structure of the contents of a code file, containing class and function declarations, without specific definitions and comments. text by syntactically analyzing the content of code files. The cross-file context is retrieved using BM25 after forming a candidate retrieval pool together with the context obtained from semantic-based retrieval. It takes into account structural information of the code but does not establish graph relations across code files. RepoFuse [64] is a solution for the Context-Latency Conundrum in repository-level code completion. It constructs Code Knowledge Graph by analyzing the code graph dependencies in the repository and uses the repository-level graphs for retrieval. It integrates the Rationale Context obtained by analyzing the repository code structure and the Analogy Context based on the retrieval of similar code blocks, and filtering the context by scoring function. RLCoder [63] is a reinforcement-learning-based framework for repository-level code completion, which can effectively improve code completion performance and has good generalization ability. During training, the RLRetriever is trained with a reward mechanism based on weighted perplexity to learn retrieval, while a stop-signal mechanism is introduced to filter candidate codes. In inference, the trained RLRetriever retrieves useful candidate codes from the code repository and inputs them together with the incomplete code into the generator to complete code generation.
@@ C.7.1 Variants of Graph RAG Framework
The Graph RAG framework, comprising Rewriter, Retriever, Reranker, and Reader, extends CGM to real-world issue fixing. In Table 10, we verify the effectiveness of each component in our Graph RAG framework by removing them. Here, we use CGM-SWE-PY (see Appendix C.1 for training details) as Reader. w/o Rewriter: We directly perform semantic search based on the original issue descriptions, obtain the anchor nodes from the code graph, and provide them to Retriever. Removing Rewriter results in an 8.33% performance drop, which proves its effectiveness in enhancing the original issue descriptions. w/o Retriever: Since there is no Retriever to provide filtered files and subgraphs, we input all the files in the original codebase into Reranker’s Stage 1 for selection, and at the same time append the key information output by Rewriter into Reranker’s prompts. Based on the files output by Reranker, we build a subgraph using these files and their one-hop neighbors, as the graph modality input of CGM. The exclusion of Retriever results in an 11.33% performance degradation, a more severe drop than removing Rewriter, highlighting its importance in providing issue-related subgraph. w/o Reranker: We use the top 5 files that are most similar to the query in embedding space (during semantic search) from the FILE node obtained by Retriever and provide them to Reader as the files to be modified. Removing Reranker results in the largest performance drop (decreased by -24.67%), emphasizing its importance in improving the precision of retrieval results and providing the right, relevant files to Reader. w/o R3: To evaluate the effectiveness of the RAG module, we create a baseline which removes the first three modules (Rewriter, Retriever, and Reranker) and feed the entire (truncated when the length exceeds the context length of the base model) repository graph as input to Reader during fine-tuning. Removing the RAG module leads to a poor performance (decreased by 33.33%), possibly due to excessive noise from the unfiltered repository graph and information loss from context-window truncation. w/o CGM Reader (FlatGraph): To verify the effectiveness of CGM Reader in jointly modeling semantics and structure, we create a naive graph-based baseline which flattens code snippets based on topological structure [24], representing an alternative Reader with structure-enhanced fine-tuning. The naive graph-based Reader only achieves 5.33% on SWE-bench Lite, far behind the proposed CGM (decreased by 37.67%).
@@ C.8 Generalization of CGM on Different Backbones
To evaluate CGM with different backbones, we trained CGM using Llama3.1-70B-Instruct, Qwen2.5-Coder-32B-Instruct, and Qwen2.5-Coder-7B-Instruct, in addition to Qwen2.5-72B. The results are summarized in Table 12. We find that the performance of CGM positively correlates with the LLM decoder’s inherent coding and instruction-following abilities. For example, Llama3.1-70B-Instruct CGM’s performance decreased 17.67% compared to Qwen2.5-72B, possibly due to weaker inherent coding abilities (see Table 2 in [20]). Still, it surpassed Lingma-SWEGPT [1] built on Llama3.1-70B-Instruct by 18.33%, demonstrating CGM’s power in improving open-source LLMs. Table 12: CGM with Different Backbones.

### v3

@@ C.1.1 Training Data
As mentioned in section 4, we construct training data for different training phase of CGM respectively. Meanwhile, to enhance the model’s ability in code completion, we also construct code completion samples for fine-tuning of code completion task. To explore the performance of CGM across different programming languages, our data includes both Python and Java. When constructing the training data, we filter out the repositories involved in the test sets for testing to avoid data leakage. Data for Subgraph Reconstruction Pre-training: We obtain 500k Python and 360k Java subgraphs (with the maximum length of 8k tokens) from a total of 20k high-star Github repositories. Data for Issue Fixing: We collect 200k issue-patch pairs (100k per language), from GitHub pull-requests. Among the 100k Python pairs, 14k are sourced from the SWEBench training set (Yang et al., 2024b). Data for Code Completion: The code completion samples are self-constructed from the above repositories, 250k per language.
@@ C.1.2 CGM-Multi
We initialize CGM-Multi with Qwen2.5-72B-Instruct (Yang et al., 2024a) as the base LLM. Then pre-train it using subgraph reconstruction data and fine-tuning data (issue-fixing and code completion) in both two languages (Python and Java). To ensure balance between different languages, we use 360k subgraph reconstruction data for both languages. Training uses 4-bit QLoRA on 64 A100s (batch=32, lr=1e-4, and epoch=2). The encoder combines CodeT5+ (Wang et al., 2023) with LoRA (rank=64, and alpha=32), and the adapter uses a two-layer MLP with GELU activation. The first layer of the adapter maps the CodeT5+ output dimension of 256 to 8192, and the second layer maintains the dimension of 8192, which aligns with the LLM’s hidden dimension. We adopt Xformers (Lefaudeux et al., 2022) for efficient attention computation.
@@ C.5.2 Datasets
The following three benchmarks, which focus on repository-level issue fixing, are all evaluated using the Docker executable environment. SWE-bench Lite (Jimenez et al., 2024): SWE-bench Lite contains 300 self-contained repository-level issues from 11 repositories, designed to test the model’s understanding of repository-level code changes and its ability to generate correct patches primarily focused on Python. It provides a realistic software engineering environment, including execution contexts, to evaluate the model’s ability to resolve real-world issues. SWE-bench Verified (OpenAI, 2024b): SWE-bench Verified contains 500 self-contained repository-level issues from 12 repositories. This dataset contains samples that have been verified to be non-problematic by human annotators. SWE-bench-java Verified (Zan et al., 2024): This dataset include 91 Java issues from 6 repositories, enabling cross-language evaluation. Like SWE-bench Lite, it provides execution environments to validate the correctness of generated patches.
@@ C.6.2 Datasets
CrossCodeEval (Ding et al., 2024) is an emerging benchmark for cross-file code completion, which is constructed from a wide range of real-world repositories from GitHub in four popular programming languages: Python, Java, TypeScript, and C#. In our experiments, we evaluate the model’s code completion ability on only two languages, Java and Python. As shown in Table 9, we provide the dataset statistics of the CrossCodeEval benchmark for Java and Python. ComplexCodeEval (Feng et al., 2024) is a new benchmark for evaluating the performance of large code models in complex development scenarios. It includes 3,897 Java samples from 1,055 Java code repositories and 7,184 Python samples from 2,107 Python code repositories. Following the original setup of this benchmark, we randomly selected 100 samples each in Python and Java for the evaluation. Table LABEL:tab:_complex_python and Table LABEL:tab:_complex_java show the information of the selected samples. When evaluating the code completion capability, unlike the original setup which requires the model to complete the second half of the function, we ask the model to complete the middle line of the function given the contextual information. Table 9: The statistics of the CrossCodeEval for Java and Python. Feature Java Python # Repositories 239 471 # Files 745 1368 # Examples 2139 2665
@@ C.6.4 Baselines: Base Model
Given the subgraph same to CGM (see Appendix C.6.1), we textualize the graph into text sequnce and input into the following baselines based on the prompt templates provided by each benchmark (Ding et al., 2024; Feng et al., 2024). Since these base models have limitations on context length, we perform truncation on text inputs larger than 8k. Mistral-Large-2 (mis, 2024) is a model developed by Mistral AI with 123 billion parameters. It stands out with a remarkable 128k tokens context length, proficiently handling dozens of languages and over 80 programming languages, excelling in code generation, mathematics, and reasoning. To be specific, we chose Mistral-Large-Instruct-2411 as the latest version of Mistral-Large to compare with CGM. DeepSeek-V2.5 (Zhu et al., 2024) is a strong Mixture-of-Experts (MoE) language model characterized by economical training and efficient inference. It comprises 236B total parameters, of which 21B are activated for each token. Qwen2.5 (Yang et al., 2024a) is a decoder-only LM series whose size varies from 0.5B to 72B trained on 18 trillion tokens. Its context length support up to 128K tokens and can generate up to 8K tokens. Qwen2.5-Coder (Hui et al., 2024) is a series of code-specific language models developed by Alibaba. Derived from Qwen2.5, it comes in six sizes, is trained on a vast 5.5-trillion-token corpus, and excels in various code-related tasks, outperforming many models of the same or larger size. Since it uses a specific format and tokens for training on the fill-in-middle code completion task, we followed this prompt setting during the evaluation to obtain its true performance. For the inference of the baseline model, we all deployed the above models using the VLLM framework with the model‘s default settings. All models inference on 4 A100s with 80G VRAM, except for DeepSeek-V2.5, which requires 8 A100s.
@@ C.6.5 Baselines: RAG Method
BM25 (Robertson et al., 2009) is a classic information-retrieval algorithm based on the probabilistic model. Its core idea is to rank documents based on the relevance between query and documents. It serves as a traditional retrieval method that does not regard the structural information naturally existing in the coding task, and only performs similarity matching based on word frequency and text length. It was used in the original CrossCodeEval dataset to search for cross-file information based on the code snippets. In our experiments, we directly use the BM25 results provided by CrossCodeEval. R2C2-Coder (Deng et al., 2024) is a method that aims to enhance and benchmark the real-world repository-level code completion abilities of code Large Language Models. In particular, $R^{2}C^{2}$-Enhance reduces cross-file information to skeleton [^2] The file skeleton is a hierarchical structure of the contents of a code file, containing class and function declarations, without specific definitions and comments. text by syntactically analyzing the content of code files. The cross-file context is retrieved using BM25 after forming a candidate retrieval pool together with the context obtained from semantic-based retrieval. It takes into account structural information of the code but does not establish graph relations across code files. RepoFuse (Liang et al., 2024) is a solution for the Context-Latency Conundrum in repository-level code completion. It constructs Code Knowledge Graph by analyzing the code graph dependencies in the repository and uses the repository-level graphs for retrieval. It integrates the Rationale Context obtained by analyzing the repository code structure and the Analogy Context based on the retrieval of similar code blocks, and filtering the context by scoring function. RLCoder (Wang et al., 2024d) is a reinforcement-learning-based framework for repository-level code completion, which can effectively improve code completion performance and has good generalization ability. During training, the RLRetriever is trained with a reward mechanism based on weighted perplexity to learn retrieval, while a stop-signal mechanism is introduced to filter candidate codes. In inference, the trained RLRetriever retrieves useful candidate codes from the code repository and inputs them together with the incomplete code into the generator to complete code generation.
@@ C.7.1 Variants of Graph RAG Framework
The Graph RAG framework, comprising Rewriter, Retriever, Reranker, and Reader, extends CGM to real-world issue fixing. In Table 10, we verify the effectiveness of each component in our Graph RAG framework by removing them. Here, we use CGM-SWE-PY (see Appendix C.1 for training details) as Reader. w/o Rewriter: We directly perform semantic search based on the original issue descriptions, obtain the anchor nodes from the code graph, and provide them to Retriever. Removing Rewriter results in an 8.33% performance drop, which proves its effectiveness in enhancing the original issue descriptions. w/o Retriever: Since there is no Retriever to provide filtered files and subgraphs, we input all the files in the original codebase into Reranker’s Stage 1 for selection, and at the same time append the key information output by Rewriter into Reranker’s prompts. Based on the files output by Reranker, we build a subgraph using these files and their one-hop neighbors, as the graph modality input of CGM. The exclusion of Retriever results in an 11.33% performance degradation, a more severe drop than removing Rewriter, highlighting its importance in providing issue-related subgraph. w/o Reranker: We use the top 5 files that are most similar to the query in embedding space (during semantic search) from the FILE node obtained by Retriever and provide them to Reader as the files to be modified. Removing Reranker results in the largest performance drop (decreased by -24.67%), emphasizing its importance in improving the precision of retrieval results and providing the right, relevant files to Reader. w/o R3: To evaluate the effectiveness of the RAG module, we create a baseline which removes the first three modules (Rewriter, Retriever, and Reranker) and feed the entire (truncated when the length exceeds the context length of the base model) repository graph as input to Reader during fine-tuning. Removing the RAG module leads to a poor performance (decreased by 33.33%), possibly due to excessive noise from the unfiltered repository graph and information loss from context-window truncation. w/o CGM Reader (FlatGraph): To verify the effectiveness of CGM Reader in jointly modeling semantics and structure, we create a naive graph-based baseline which flattens code snippets based on topological structure (Guo et al., 2024), representing an alternative Reader with structure-enhanced fine-tuning. The naive graph-based Reader only achieves 5.33% on SWE-bench Lite, far behind the proposed CGM (decreased by 37.67%).
@@ C.8 Generalization of CGM on Different Backbones
To evaluate CGM with different backbones, we trained CGM using Llama3.1-70B-Instruct, Qwen2.5-Coder-32B-Instruct, and Qwen2.5-Coder-7B-Instruct, in addition to Qwen2.5-72B. The results are summarized in Table 12. We find that the performance of CGM positively correlates with the LLM decoder’s inherent coding and instruction-following abilities. For example, Llama3.1-70B-Instruct CGM’s performance decreased 17.67% compared to Qwen2.5-72B, possibly due to weaker inherent coding abilities (see Table 2 in (Yang et al., 2024a)). Still, it surpassed Lingma-SWEGPT (Ma et al., 2024a) built on Llama3.1-70B-Instruct by 18.33%, demonstrating CGM’s power in improving open-source LLMs. Table 12: CGM with Different Backbones.